A firm employs 300 women and 100 men. The mean number of days absent last year for the women was 5.3 with a s.d. of 2.2 and for the men the corresponding figures were 6.2 and 2.9. Is the difference between the means significant at 5% level of significance?


1
Expert's answer
2022-02-25T05:53:36-0500

A F-test is used to test for the equality of variances. The following F-ratio is obtained: "F=\dfrac{s_1^2}{s_2^2}=\dfrac{2.2^2}{2.9^2}\approx0.5755"

The critical values for "df_1=300-1=299" degrees of freedom, "df_2=100-1=99" degrees of freedom, are "F_L = 0.7337" and "F_U = 1.3995," and since "F = 0.5755," then the null hypothesis of equal variances is rejected.

The following null and alternative hypotheses need to be tested:

"H_0:\mu_1=\mu_2"

"H_1:\mu_1\not=\mu_2"

This corresponds to a two-tailed test, for which a t-test for two population means, with two independent samples, with unknown population standard deviations will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the degrees of freedom are computed as follows, assuming that the population variances are unequal:


"df_{total}=\dfrac{(\dfrac{s_1^2}{n_1}+\dfrac{s_2^2}{n_2})^2}{\dfrac{(s_1^2/n_1)^2}{n_1-1}+\dfrac{(s_2^2/n_2)^2}{n_2-1}}"

"=\dfrac{(\dfrac{2.2^2}{300}+\dfrac{2.9^2}{100})^2}{\dfrac{(2.2^2/300)^2}{300-1}+\dfrac{(2.9^2/100)^2}{100-1}}"

"=138.9337"

Hence, it is found that the critical value for this two-tailed test is "t_c = 1.9772," for "\alpha = 0.05"  and "df=138.9337."

The rejection region for this two-tailed test is "R = \{t: |t| > 1.9772\}."

Since it is assumed that the population variances are unequal, the t-statistic is computed as follows:


"t=\dfrac{\bar{X}_1-\bar{X}_2}{\sqrt{\dfrac{s_1^2}{n_1}+\dfrac{s_2^2}{n_2}}}"

"=\dfrac{5.2-6.3}{\sqrt{\dfrac{2.2^2}{300}+\dfrac{2.9^2}{100}}}=-2.8427"

Since it is observed that "|t| = 2.8427 > t_c = 1.9772," it is then concluded that the null hypothesis is rejected.

Using the P-value approach: The p-value for "df=138.9337, t=-2.8427," two-tailed is "p=0.005148," and since "p=0.005148<0.05=\alpha," it is concluded that the null hypothes is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu_1" is different than "\mu_2," at the "\alpha = 0.05" significance level.